Nesmith Corporation's outstanding bonds have a $1,000 par value, an 11% semiannual coupon, 12 years to maturity, and an 8% YTM. What is the bond's price? Round your answer to the nearest cent.
Answer #1

Price of a bond is mathematically represented as: 1 1-- (1 + i) PEC. - +- м. (1 + i) where P is price of bond, with periodi

M = $1000, n = 12 * 2 = 24 semi-annual periods, i = 8%/2 = 4% (semi-annual), C = 11% * $1000/2 = $55

1- P = 55* (1+0.04 24 - 0.04 + 1000 (1 + 0.04)24

P = $838.58 + $390.12

P = $1,228.7